# Learning What to hear: Boosting Sound-Source Association for Robust Audiovisual Instance Segmentation

### v1

@@ Abstract
Audiovisual instance segmentation (AVIS) requires accurately localizing and tracking sounding objects throughout video sequences. Existing methods suffer from visual bias stemming from two fundamental issues: uniform additive fusion prevents queries from specializing to different sound sources, while visual-only training objectives allow queries to converge to arbitrary salient objects. We propose Audio-Centric Query Generation using cross-attention, enabling each query to selectively attend to distinct sound sources and carry sound-specific priors into visual decoding. Additionally, we introduce Sound-Aware Ordinal Counting (SAOC) loss that explicitly supervises sounding object numbers through ordinal regression with monotonic consistency constraints, preventing visual-only convergence during training. Experiments on AVISeg benchmark demonstrate consistent improvements: +1.64 mAP, +0.6 HOTA, and +2.06 FSLA, validating that query specialization and explicit counting supervision are crucial for accurate audiovisual instance segmentation. Our code and models are available at https://github.com/jinbae-s/ACVIS
Index Terms—  Audio-visual instance segmentation, Multimodal Learning, Ordinal regression
@@ 1 Introduction
Early audiovisual learning explored multimodal correspondence through self-supervision [2], evolving from coarse localization [3] to attention-based [4] and contrastive methods [5]. However, these methods are limited to producing only heat maps or bounding boxes rather than semantic-level masks. To address this limitation, audiovisual segmentation emerged with pixel-level precision, starting with AVS [6, 7] that introduced temporal pixel-wise audiovisual interaction. Subsequent works—AVSegFormer [8] with transformers, SAMA-AVS [9] leverages SAM [10], and VCT [11] with vision-centric queries—improved semantic segmentation but all remained limited to semantic-level, unable to distinguish individual instances. Recently, AVISM [1] has achieved instance-level segmentation through a compacted two-stage architecture: frame-level object localization followed by video-level object tracking. In the first stage, the model predicts frame queries corresponding to instances for each frame by adding audio features to learnable query tokens, which are then refined through visual cross-attention to produce audiovisual frame queries. In the second stage, these frame-level detections are associated across time through a tracker that establishes temporal correspondences between instances. While this approach enables instance-level tracking, the frame-level localizer suffers from a critical limitation: audio features are uniformly integrated into all queries through a simple addition process, preventing queries from specializing to different sound sources.
@@
  • We introduce a novel Audio-Centric audioVisual Instance Segmentation (ACVIS) for sound source-aware audiovisual instance segmentation, introducing an audio-centric query generator with sound-aware ordinal counting loss.
@@
  • We demonstrate through experiments where improvements of +1.64 mAP, +0.6 HOTA, and +2.06 FSLA, validating our frame-level innovations as crucial for accurate audio-visual instance segmentation.
@@ 2.1 Problem Formulation and Overview
Audiovisual instance segmentation (AVIS) aims to classify, segment, and track all sounding objects in a given video. Given an input video with audio, the model predicts a set of instance masks with associated class labels.
Our method takes AVISM [1] as a baseline model, which follows the set-prediction paradigm [12, 13] with a two-stage architecture: frame-level object localizer and video-level object tracker. Formally, audio and video encoders extract audio feature $f^{A}_{t}$ and visual feature $f^{V}_{t}$ for a given input, respectively. The pixel decoder within the object localizer generates enhanced multi-scale features: $\bar{f}^{V}_{t}$ for final-resolution map and $\bar{\mathcal{F}}^{V}_{t}$ for multi-scale representations. These features, combined with audio features $f^{A}_{t}$ and learnable queries $\mathbf{q}$, produce audiovisual frame queries $\mathbf{q}^{AV}_{t}$ at frame $t$. The object tracker aggregates these frame queries $\{\mathbf{q}^{AV}_{t}\}_{t=1}^{T}$ into $N_{v}$ video queries to generate final predictions $\hat{\mathcal{Y}}$:
@@ 2.3 Sound-Aware Ordinal Counting loss
To enhance the sound-source awareness in our framework, we design a count token $q_{\text{cnt}}\in\mathbb{R}^{D}$ to optimize the weights with our SAOC loss. The count token is concatenated with frame queries, and aggregates the information about the number of sound sources in the object localizer. Following previous work [14], we model counting sound instances as an ordinal regression problem using conditional probabilities to ensure rank consistency. The count token $q_{\text{cnt}}^{AV}$ is processed by $\phi_{\text{cnt}}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{K_{\max}}$ to predict:
@@ 2.4 Training and Inference
where $\lambda_{\text{SAOC}}$ is the hyperparameter for $\mathcal{L}_{\text{SAOC}}$ and $\mathcal{L}_{\text{AVIS}}$ is the weighted sum of $\mathcal{L}_{\text{frame}}$, $\mathcal{L}_{\text{video}}$ and $\mathcal{L}_{\text{sim}}$.
Inference. During inference, the $N_{v}$ video-level predictions are filtered through confidence thresholding to produce $N_{\text{pred}}$ final instance trajectories $\hat{\mathcal{Y}}=\{(\hat{m}_{n},\hat{c}_{n})\}_{n=1}^{N_{\text{pred}}}$, where each instance spans multiple frames with mask $\hat{m}_{n}\in[0,1]^{T\times H\times W}$ and class logits $\hat{c}_{n}\in\mathbb{R}^{K+1}$.
@@ 3 Experiments
Model mAP HOTA FSLA FSLAn FSLAs FSLAm
AVISM 45.04 64.52 44.42 20.62 32.62 54.99
@@
Table 1: Performance comparisons with SoTAs.
@@
Table 3: Impact of audio-centric query generator and $\mathcal{L}_{\text{SAOC}}$
@@
Table 5: Impact of $K_{\text{max}}$
@@ 4 Conclusion
We presented audio-centric query generation and sound-aware ordinal counting loss to address visual dominant regime in audiovisual instance segmentation. Our method enables queries to specialize to different sound sources, while SAOC loss explicitly supervises sounding object counts via ordinal regression. Comprehensive experiments demonstrate that both query specialization and counting supervision are essential, achieving strongest gains in challenging multi-source scenario. These results validate that balanced frame-level processing via both architectural and objective design is crucial for accurate audiovisual instance segmentation.

### v2

@@ Abstract
Audiovisual instance segmentation (AVIS) aims to accurately localize and track sounding objects throughout video sequences. Existing methods suffer from visual bias stemming from two fundamental issues: uniform additive fusion prevents queries from specializing to different sound sources, while visual-only training objectives limit queries from converging to arbitrary salient objects. We propose Audio-Centric Query Generation (ACQG) using cross-attention mechanism, enabling each query to selectively attend to distinct sound sources and carry sound-specific priors into visual decoding. Additionally, we introduce Sound-Aware Ordinal Counting (SAOC) loss that explicitly supervises sounding object numbers through ordinal regression with monotonic consistency constraints, preventing visual-only convergence during training. Experiments on AVISeg benchmark demonstrate consistent improvements: +1.64 mAP, +0.6 HOTA, and +2.06 FSLA, validating that query specialization and explicit counting supervision are crucial for accurate audiovisual instance segmentation. Our code and models are available at https://github.com/jinbae-s/ACVIS
Index Terms—  Audiovisual instance segmentation, Multimodal Learning, Ordinal regression
@@ 1 Introduction
Early audiovisual learning explored multimodal correspondence through self-supervision [1], evolving from coarse localization [18] to attention-based [17] and contrastive methods [3]. However, these methods are limited to producing only heat maps or bounding boxes rather than semantic-level masks. To address this limitation, audiovisual segmentation [22, 21] has emerged with pixel-level precision. Subsequent works—AVSegFormer [6] with transformers, SAMA-AVS [14] leverages SAM [11], and VCT [10] with vision-centric queries—have improved semantic segmentation but all remained limited to semantic-level, unable to distinguish individual instances. Recently, AVISM [7] has achieved instance-level segmentation through a compacted two-stage architecture: frame-level object localization followed by video-level object tracking. In the first stage, the model predicts frame queries corresponding to instances for each frame by adding audio features to learnable query tokens, which are then refined through visual cross-attention to produce audiovisual frame queries. In the second stage, these frame-level detections are associated across time through a tracker that establishes temporal correspondences between instances. While this approach enables instance-level tracking, the frame-level localizer suffers from a critical limitation: audio features are uniformly integrated into all queries through a simple addition process, preventing queries from specializing to different sound sources.
@@
  • We introduce a novel Audio-Centric audioVisual Instance Segmentation (ACVIS) for sound source-aware AVIS, introducing an audio-centric query generator (ACQG) with sound-aware ordinal counting (SAOC) loss.
@@
  • We demonstrate through experiments where improvements of +1.64 mAP, +0.6 HOTA, and +2.06 FSLA, validating our frame-level innovations as crucial for accurate AVIS.
@@ 2.1 Problem Formulation and Overview
AVIS aims to classify, segment, and track all sounding objects in a given video. Given an input video with audio, the model predicts a set of instance masks with associated class labels.
Our method takes AVISM [7] as a baseline, which follows the set-prediction paradigm [2, 4] with a two-stage architecture: frame-level object localizer and video-level object tracker. Formally, audio and video encoders extract audio feature $f^{A}_{t}$ and visual feature $f^{V}_{t}$ for a given $t$-th segment, respectively. The pixel decoder within the object localizer generates enhanced multi-scale features: $\bar{f}^{V}_{t}$ for final-resolution map and $\bar{\mathcal{F}}^{V}_{t}$ for multi-scale representations. These features, combined with audio features $f^{A}_{t}$ and learnable queries $\mathbf{q}$, produce audiovisual frame queries $\mathbf{q}^{AV}_{t}$ at frame $t$. The object tracker aggregates these frame queries $\{\mathbf{q}^{AV}_{t}\}_{t=1}^{T}$ into $N_{v}$ video queries to generate final predictions $\hat{\mathcal{Y}}$:
@@ 2.3 Sound-Aware Ordinal Counting loss
To enhance the sound-source awareness in our framework, we design a count token $q_{\text{cnt}}\in\mathbb{R}^{D}$ to optimize the weights with our SAOC loss. The count token is a learnable embedding that is concatenated with frame queries, and aggregates the information about the number of sound sources in the object localizer. We denote by $q^{AV}_{cnt}$ the count token after the segmentation decoder. Following previous work [19], we model counting sound instances as an ordinal regression problem using conditional probabilities to ensure rank consistency. The count token $q_{\text{cnt}}^{AV}$ is processed by linear projection head $\phi_{\text{cnt}}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{K_{\max}}$ to predict:
@@ 2.4 Training and Inference
where $\lambda_{\text{SAOC}}$ is the hyperparameter for $\mathcal{L}_{\text{SAOC}}$, and $\mathcal{L}_{\text{AVIS}}$ is the weighted sum of $\mathcal{L}_{\text{frame}}$, $\mathcal{L}_{\text{video}}$ and $\mathcal{L}_{\text{sim}}$, as defined in AVISM [7].
Inference. During inference, the $N_{v}$ video-level predictions are filtered through confidence thresholding to produce $N_{\text{pred}}$ final instance trajectories $\hat{\mathcal{Y}}=\{(\hat{m}_{n},\hat{c}_{n})\}_{n=1}^{N_{\text{pred}}}$, where each instance spans multiple frames with mask $\hat{m}_{n}\in[0,1]^{T\times H\times W}$ and class logits $\hat{c}_{n}\in\mathbb{R}^{N_{c}+1}$, where $N_{c}$ is the number of object classes.
@@ 3 Experiments
Method mAP HOTA FSLA FSLAn FSLAs FSLAm
AVISM [7] 45.04 64.52 44.42 20.62 32.62 54.99
@@
Table 1: Performance comparisons on AVISeg.
@@
Table 3: Impact of audio-centric query generator and $\mathcal{L}_{\text{SAOC}}$.
@@
Table 5: Impact of $K_{\text{max}}$.
@@ 4 Conclusion
We propose audio-centric queries and sound-aware ordinal counting loss to address visual dominance in AVIS. By specializing queries and supervising counts, our method achieves significant gains in multi-source scenarios, validating the importance of balanced frame-level processing.
Acknowledgement. This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2025-02216328).